Sharon is 42 years old. Her daughter is 12 years old. How many years ago was Sharon four times older than her daughter?
Answer by lwsshak3(11628)

Sharon is 42 years old. Her daughter is 12 years old. How many years ago was Sharon four times older than her daughter?
let x= years ago when Sharon four times older than her daughter?
(42-x)=sharon's age x yrs ago
(12-x)=daughter's age x yrs ago
42-x=4(12-x)
42-x=48-4x
3x=6
x=2
How many years ago was Sharon four times older than her daughter? 2